-1/6x(2x+12)=-1/3x+12 solve for X
I no it's no solution, but I need to see it worked out to understand.

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

-1/6x(2x+12)=-1/3x+12 solve for X
--------------
It's ambiguous. Add parentheses.

Answer by ikleyn(52785)

.
-1/6x(2x+12)=-1/3x+12 solve for X
I no it's no solution, but I need to see it worked out to understand.
~~~~~~~~~~~~~~~~~~~

To simplify, multiply both sides by 6.  You will get

    -x*(2x+12) = -2x + 72.


Simplify further

    -2x^2 - 12x = -2x + 72.


Collect all the term in one side

    -2x^2 - 10x - 72 = 0.


Change the sign everythere

    2x^2 + 10x + 72 = 0.


Divide by 2

    x^2 + 5x + 36 = 0


Compute the discriminant of this quadratic equation

    d = b^2 - 4ac = 5^2 - 4*1*36.


Without calculations, it is clear that the discriminant is negative, 
which means that the equation HAS NO real solution/solutions.


It is what they want to be proven.

Solved, with complete explanations.

------------------

If you want to see it visually, go to web-site

www.desmos.com/calculator

and use free of charge plotting tool there.

Print your equations in the specialized windows, and you will see
that the parabola of the left side and the straight line of the right side
do not have common intersection points.

It is the precise geometric meaning of the problem' statement.